OpenStudy (anonymous):

integrate sin^4xcos^5xdx

OpenStudy (anonymous):

Note that: ∫ sin^3(x) * cos^4(x) dx = ∫ sin^2(x) * cos^4(x) [sin(x) dx] = ∫ [1 - cos^2(x)] * cos^4(x) [sin(x) dx] = - ∫ [1 - cos^2(x)] * cos^4(x) [-sin(x) dx]. Then, letting u = cos(x) <==> du = -sin(x) dx yields: - ∫ [1 - cos^2(x)] * cos^4(x) [-sin(x) dx] = - ∫ (1 - u^2) * u^4 du = - ∫ u^4 - u^6 du = ∫ u^6 - u^4 du = ∫ u^6 du - ∫ u^4 du = (1/7)u^7 - (1/5)u^5 + C = (1/7)cos^7(x) - (1/5)cos^5(x) + C. <== ANSWER I hope this helps!

OpenStudy (anonymous):

hihi

OpenStudy (anonymous):

he looks right

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

@Rohangrr what happen to the other sin x and cosx

OpenStudy (anonymous):

still there? have any questions?

OpenStudy (anonymous):

yes im still here I was just confused where was the other sinx cosx

OpenStudy (anonymous):

looking at rohangff work, what line?

OpenStudy (anonymous):

in that first line where sin^(3)x cos^(4)x dx did he factor that out then what happen to the other one.?

OpenStudy (anonymous):

hmm, you're right... i don't think he did your problem... double checking...

OpenStudy (anonymous):

I really dont know how to integrate this what formula did he used?

OpenStudy (anonymous):

yea, he didn't do the problem but the method should be the same...

OpenStudy (anonymous):

ok, lemme try..

OpenStudy (anonymous):

thank you so much then

OpenStudy (anonymous):

 you follow up to that last line?